# Key Words.: binaries: spectroscopic – binaries: close – stars: early-type – stars: massive – Magellanic Clouds1 Introduction

### v1

@@ 4.1 Variations of the observed multiplicity fraction
4.2 Intrinsic multiplicity fraction
We then applied our binary detection criteria described in Sect. 3.2 to the simulated observations to identify which synthetic binaries would be detected in our survey. By repeating this process 10 000 times, we obtained robust statistics and estimated the detection probability () as the fraction of synthetic binaries that met our detection criteria. This allowed us to compute the expected observed multiplicity fraction () given and intrinsic multiplicity fraction .
@@ 4.2 Intrinsic multiplicity fraction
The detected probabilities for our sample of B-type binaries are shown in Fig. 9. As expected, we have a very high detection probability for short-period systems, for orbital periods up to 10 d. After 60-70 d the detection probability drops significantly, reaching for d and for d, with an integrated probability for orbital periods between 10 d and 1 yr of . When considering the full possible orbital periods range () the detection probability of the BLOeM campaign for the B-type stars is . In the case of the mass ratio, the sharp drop is due to the minimum value considered (see Table 3). None of , , or have as much effect in the detection probability as .
After accounting for the observational biases in the BLOeM campaign and given our detection probability, we found an intrinsic multiplicity fraction of . We can compare this value to the intrinsic multiplicity fraction found for the B-type stars in 30 Dor of (Dunstall et al., 2015). Performing a Z-test, we found these two multiplicity fractions to deviate by , which is statistically significant at the 1% –level (p=0.0015). The higher SMC fraction is even more relevant if we consider that the 30 Dor sample was dominated by luminosity class V stars (close to 60%) whereas our III/V sample contains a majority of giant stars, which as seen in Fig. 5, cannot avoid interactions below orbital periods of a few days. The difference is even larger () when compared to of the galactic B-type stars in NGC 6231 (Banyard et al., 2022), and to (, assuming a mean uncertainty of 0.10) found for the OB-type stars in the Cygnus OB2 association (Kobulnicky et al., 2014). This suggests a trend with metallicity that will be explored in more detail in Sect. 6.1.
@@ 5.1 Eclipsing binaries
Some of our targets have been identified as eclipsing binaries (EB) by OGLE (Pawlak et al., 2016). We found 41 EBs which are listed in Table 5, including their photometric periods, -band magnitudes, and light curve type. Their distribution of -band magnitudes and orbital periods are shown in Fig. 10. The 41 EBs correspond to 27% of our binary sample. Among them, there are 10 SB1 systems, 26 SB2s, all three SB3s, and two RV var stars. The latter two systems did not meet our binary criteria, but we found the correct orbital period for one (BLOeM 6-061), and the second system (BLOeM 7-054) has an orbital period just outside our frequency grid (1.099 d, see Sect. 5.2).
Most of the eclipsing binaries have been classified as detached or semi-detached by Pawlak et al. (2016), but there are two ellipsoidals and one contact system. The ellipsoidal systems are both SB2s that suffered from misidentification of their components, however, we still found half of the photometric orbital period in both cases, as is common for ellipsoidal variables. The only contact system, BLOeM 6-010, also suffered from component misidentification and it was flagged as a possible SB3, but we found the correct orbital period of 1.21 d. Despite finding the same period (or half in some cases) as the OGLE period for these systems, in the three cases the significance of the period found was too low for them to have been listed as the correct periods (see details below).
We have done a preliminary inspection of the photometric data of the EBs mainly to find evidence and confirm third components. As mentioned in Sec. 3.2, we found clear signs of third companions in three systems that we have classified as SB3, but also evidence of potential contributions from further components in other 15 systems. Our photometric analysis indicates that there are at least eight systems where there is likely a third companion, which would represent a lower limit of 3% of triple systems in our sample. We provide comments on these and other systems in Appendix C.
@@ 5.3.1 Orbital periods
The B-type Binaries Characterisation programme (BBC; Villaseñor et al., 2021) studied 88 of the binary candidates identified by Dunstall et al. (2015) with 29 FLAMES observations over a baseline of 427 d. Villaseñor et al. (2021) found orbital solutions for 50 SB1 and 14 SB2 systems, and potential solutions for 20 more systems. We can compare the distributions of orbital periods for both populations, over a similar baseline. For this, we have selected 9 epochs from the BBC campaign covering a baseline of 60 d, without including the low S/N epochs 10, 12, 13, 17 (see their Table 1 and data reduction section). We computed the LS periodogram for the corresponding RVs for all systems in the BBC sample.
The results can be seen in Fig. 12. Similarly as with the periods for the BLOeM sample, we distinguish between those systems whose PLS exceeds the 0.1% FAL (turquoise circles, 29 system), those between the 0.1 and 1% thresholds (orange rectangles, 13 systems), and those with PLS below the 1% threshold (purple diamonds, 45 systems). We found excellent agreement for the first two groups. When limiting the results to a maximum period of 60 d, we found 93% agreement for systems with PLS above the 0.1% threshold, and 90% agreement for those between 0.1–1%. For this reason we use the periods from both these groups in the comparison with the BLOeM sample, resulting in a sample of 34 systems.
In the case of the BLOeM periods, our test with the BBC sample reaffirm our confidence in the periods that met our significance criterion. However, we restrict the periods to those highlighted in Fig. 11, not including those with P FAL. We do include the orbital periods from OGLE, for a total of 57 objects.
The cumulative distribution of orbital periods is shown in Fig. 13. For comparison, we also plotted the distribution of periods for all BBC systems with a period below 60 d (green dashed line). When comparing both samples with nine epochs, there is only a marginal difference at 6 d. To assess the significance of this difference we have computed the Kolmogorov-Smirnov (KS) test, obtaining a KS-statistic of and a p-value of 0.16, which does not reach significance at the 10% level. However, when comparing with the full BBC sample up to periods of 60 d, we do find significant differences (, p=0.0003). It is clear that both samples with nine epochs lack system with periods larger than d in contrast to the full BBC sample which approaches a uniform distribution in . Judging from Fig. 12, the significance of the periods found with the LS periodogram starts to decrease after 10 d, and from that point up to 60 d we see a majority of systems with P FAL, even though we did find the correct periods for many of them.
@@ 5.3.2 Distribution of
We also examined the distribution of values for the BBC systems with determined periods. Given the larger number of epochs (29) of the BBC programme and the reliable periods, their values are robust. The distribution is shown in Fig. 14, where we have colour-coded the stacked histogram by the binary classifications used by Villaseñor et al. (2021). The top panel shows the values computed from all the epochs with well determined RV measurements that were used in the computation of orbital solutions. In the bottom panel, we plot the distribution obtained by only considering nine epochs as in the case of the BLOeM sample. Comparing the two panels, we see that the distribution is shifted slightly towards lower values when only nine epochs are used, with long-period, eccentric binaries being the most affected. Indeed, the three SB1 systems that fall below 20 km s-1 all have d. Notably, 34 of the BBC SB1 systems have values below 80 km s-1, using a large binary threshold such as this (see discussion in Sect. 6.2) would miss 64% of the SB1 systems. On the other side, a threshold of 20 km s-1 would still skip some of the confirmed long-period binaries in BBC when only using nine epochs, whereas only three systems without orbital solutions remain above the threshold. This evidence also suggest that 20 km s-1 is an appropriate limit to separate binarity from intrinsic variability.
@@ 6.1 Trend with metallicity
As seen in Sect. 4.2, the large multiplicity fraction we have found for the BLOeM B-type stars suggests a trend with metallicity. To test this dependence, we performed a simple linear regression (as the one applied by Sana et al. 2024 to the O-type stars) to the results from Kobulnicky et al. (2014); Dunstall et al. (2015); Banyard et al. (2022), and our result, where we have adopted as uncertainty for the former the mean of the errors of the other studies, and for simplicity, in the case of asymmetric errors, the mean between them. The fit is shown in Fig. 15, with its 68% confidence interval in shaded blue. We found a slope of , which is away from 0.
To illustrate and compare with the case of solar-type stars, we include in the plot a sample of G- and K-type IV/V stars observed by APOGEE and first presented by Badenes et al. (2018). Close binary fraction values as a function of metallicity were obtained from Moe et al. (2019, their figure 11), with the corresponding fit shown in red. We also include three additional values (red diamonds, not included in the linear regression) at [Fe/H]1.0, 0.2, and 0.5. These points represent a weighted moving average from five different solar-type samples, including the Badenes et al. sample, analysed by Moe et al. (2019, see their figure 18). The averaged values are in good agreement with the former measurements, but we only fit the APOGEE sample since it is much larger ( objects) and presents the lowest errors. For the details of the solar-type samples we refer the reader to Moe et al. (2019), but we note that all samples were corrected for incompleteness up to orbital periods of . We have computed the linear fit to the APOGEE data and found a slope of . The two slope values for solar and B-type stars are not significantly different from each other when compared with a t-test (), suggesting a similar trend of the multiplicity fraction with metallicity. Figure 15 also shows the linear fit found by Sana et al. (2024) for the O-type stars (purple dashed line) with its corresponding MW (Sana et al., 2012), LMC (Sana et al., 2013; Almeida et al., 2017), and SMC (Sana et al., 2024) multiplicity fractions. No clear relation was found for the O-type stars with metallicity by Sana et al. (2024).
@@
Close massive binaries ( au) are believed to form through disk fragmentation followed by inward orbital migration via circumbinary accretion (Tokovinin & Moe, 2020). Those in the closest orbits ( d) require extremely early fragmentation, coupled with subsequent envelope mass accretion to harden the orbits to such short periods, a process that remains largely independent of metallicity. This would explain why EBs, which predominantly probe these very close binaries, exhibit multiplicity properties that are invariant with metallicity. This is similar to what we see for the more massive O-type stars. In order to form an O-type star, a high accretion rate and large disk surface density are needed, which do not depend on metallicity. This presents a natural explanation for why the multiplicity fraction of O-type stars remains high at different metallicities.
However, for B-type binaries at wider separations ( au), metallicity might have played a more significant role. Metal-poor accretion disks, such as those in the SMC, are less optically thick and can cool more efficiently. This enhanced cooling and reduced opacity facilitate gravitational instabilities, leading to increased disk fragmentation. Therefore, for stars where fragmentation occurs at later times, there is not enough remaining disk mass or accretion to dissipate the orbital energy effectively and decrease separation beyond au. This mechanism would potentially result in a larger multiplicity fraction in early B-type binaries with separations between 0.2 and 10 au at low metallicity.
@@ 6.2 Considering a higher RV threshold
An alternative explanation is that our assumed distributions of orbital parameters are incorrect. Our results are particularly sensitive to the orbital period distribution, for which we have assumed an Öpik law (see Table 3). As discussed in Sect. 5, the fraction of binary system with periods between –10 d seems to be larger for the BLOeM sample than for the BBC sample of binaries. Although our analysis of the orbital period distribution is inconclusive about the significance of the difference in the distribution of both samples according to the KS test, there is other supporting evidence that is worth examining. First, the fraction of SB2 systems is much larger among the BLOeM B-type binaries in comparison to the 30 Dor counterpart, with 39% of the binaries (41% if we include the SB3 systems) versus only 17% in the BBC sample. There are several factors influencing the detection of SB2 systems, some are intrinsic to the system such as orbital period, mass ratio, and rotational broadening, but external factors like S/N can also play a role. Villaseñor et al. (2021) estimated a mean S/N for the BBC sample of 33. In contrast, the BLOeM sample has a mean S/N of 55, which is significantly higher, probably responding to the higher fraction of giant stars. Moreover, the fraction of SB2s is similar to that of the VFTS O-type binaries (Almeida et al., 2017), which presented similar S/N to our sample. In conclusion, the fraction of SB2 systems, although large, is not impossible to explain.
EBs are also abundant among the BLOeM B-type III/V sample, reaching 27% of the binaries, compared to only 12% found by the BBC programme (Villaseñor et al., 2021). This could also be seen as evidence of an overabundant fraction of short-period systems. However, contrarily to BBC, our BLOeM sample is dominated by giant stars which are roughly a factor two larger than their MS counterpart. Since the probability of eclipses scales with radius, giant primaries will have a higher probability of being seen as EBs. Only accounting for MS stars, we find that similar fractions of the BBC and BLOeM binaries are MS EBs (14% and 11% respectively).
@@ 6.3 Binary fraction from the distribution of
We adopt an MC method similar to that described in Sect. 4.2, sampling the same physical and orbital parameters but extending the orbital period range down to to include sub-day periods observed by OGLE. We also extend the mass-ratio distribution to to probe lower-mass companions. We assumed circular orbits () for systems with orbital periods below 2 d (Moe & Di Stefano, 2017), whereas for longer orbits we sample from 0 to (as given by Badenes et al. (2018)). Values of the distribution indices remain as in Table 3. To reduce stochastic noise, we generate 10 000 stars with a binary fraction given by and use a MCMC approach to fit the observed distribution by varying . We scaled each simulated distribution by the ratio of observed to simulated stars, and then compare with the real data using a Poisson-based log-likelihood in log-space intervals:
@@
The posterior distribution for is shown in Fig 16, our simulations clearly favour a high binary fraction of % in good agreement with our value of % reported in Sect. 4.2. Figure 17 shows the observed distribution (black histogram), it can be seen that there is no evident separation between a core and tail component, possibly due to the uncertainty in our RV measurements or the scarcity of very wide binaries () in comparison to the distribution of orbital periods of sun-like stars which peaks roughly at d (see Raghavan et al., 2010; Badenes et al., 2018). We include the distribution of values for four simulated samples of 10 000 stars computed with an intrinsic binary fraction of 0.0, 0.5, 0.78, and 1.0 to illustrate the change in the distribution with . A lower binary fraction naturally increases the number of stars with low-amplitude RV variations, whereas a higher binary fraction adds more highly RV-variable systems. Notably, the transition value between single- and binary-dominated bins in the simulations occurs near 20 km s-1, supporting the choice of this value as a sensible threshold.
Another point worth mentioning is the inclusion of slightly evolved stars in our sample. As shown in Sect. 4 and Fig. 5, for typical and detached systems with evolved primaries () can exhibit peak-to-peak RVs about 100 km s-1 lower than those with . These findings illustrate how stellar evolution can effectively shift some of the high-amplitude binaries to lower values in the observed distribution of . Because our simulations do not track orbital evolution with mass transfer or stellar ageing, these post-interaction systems might scape detection. Future BLOeM papers will explore these evolutionary effects in detail. In conclusion, our multiplicity fraction can thus be interpreted as a lower limit due to binary evolution likely removing part of the high- tail from the distribution.
It is also interesting to examine the distribution of values for (blue histogram in Fig. 17); there is a fraction of simulated single stars (9%) with values above 20 km s-1. The RV values are simulated by adding a Gaussian uncertainty with a standard deviation corresponding to the uncertainty of the RV value at each epoch of a real, randomly assigned, BLOeM star. One single bad epoch (e.g. large uncertainty due to low S/N or a bad multi-component fit) can produce a large value. Figure 18 shows the distribution of the simulated sample of single stars in the binary criteria diagram. Out of the 10 000 stars, only four pass both binary criteria, meaning only a 0.04% of false-positive detections. This exercise not only confirms the high multiplicity fraction of early B-type dwarfs and giants in the SMC, but also further validates the choice of the binary criteria for our sample.
6.4 Variations of the orbital period distribution

### v4

@@ 4.3 Intrinsic multiplicity fraction
We then applied our binary detection criteria described in Sect. 3.2 to the simulated observations to identify which synthetic binaries would be detected in our survey. By repeating this process 10 000 times, we obtained robust statistics and estimated the detection probability () as the fraction of synthetic binaries that met our detection criteria. This allowed us to compute the expected observed multiplicity fraction () given an intrinsic multiplicity fraction .
@@
The detection probabilities for our sample of B-type binaries are shown in Fig. 8. As expected, we have a very high detection probability for short-period systems, for orbital periods up to 10 d. After 60-70 d the detection probability drops significantly, reaching for d and for d, with an integrated probability for orbital periods between 10 d and 1 yr of . When considering the full possible orbital periods range () the detection probability of the BLOeM campaign for the B-type stars is . In the case of the mass ratio, the sharp drop is due to the minimum value considered (see Table 3). None of , , or have as much effect in the detection probability as .
After accounting for the observational biases in the BLOeM campaign and given our detection probability, we found an intrinsic multiplicity fraction of . We can compare this value to the intrinsic multiplicity fraction found for the B-type stars in 30 Dor of (Dunstall et al., 2015). Performing a Z-test, we found these two multiplicity fractions to deviate by , which is statistically significant at the 1% –level (p=0.0015). The higher SMC fraction is even more relevant if we consider that the 30 Dor sample was dominated by luminosity class V stars (close to 60%) whereas our III/V sample contains a majority of giant stars, which as seen in Fig. 4, cannot avoid interactions below orbital periods of a few days. The difference is even larger () when compared to of the galactic B-type stars in NGC 6231 (Banyard et al., 2022), and to (, assuming a mean uncertainty of 10) found for the OB-type stars in the Cygnus OB2 association (Kobulnicky et al., 2014). This suggests a trend with metallicity that will be explored in more detail in Sect. 6.1.
@@ 5.1 Eclipsing binaries
Some of our targets have been identified as EBs by OGLE (Pawlak et al., 2016). We found 41 EBs which are listed in Table D.2, including their photometric periods, -band magnitudes, and light curve type. Their distribution of -band magnitudes and orbital periods are shown in Fig. 9. The 41 EBs correspond to 27% of our binary sample. Among them, there are ten SB1 systems, 26 SB2s, all three SB3s, and two RV var stars. The latter two systems did not meet our binary criteria, but we found the correct orbital period for one (BLOeM 6-061), and the second system (BLOeM 7-054) has an orbital period just outside our frequency grid (1.099 d, see Sect. 5.2).
Most of the EBs have been classified as detached or semi-detached by Pawlak et al. (2016), but there are two ellipsoidals and one contact system. Both ellipsoidal systems are SB2s that suffered from misidentification of their components. However, we still recovered half of the photometric orbital period in both cases, as is common for ellipsoidal variables. The only contact system, BLOeM 6-010, also suffered from component misidentification and it was flagged as a possible SB3, but we found the correct orbital period of 1.21 d. Despite finding the same period (or half in some cases) as the OGLE period for these systems, the significance in all three cases was too low to confidently list them as the correct periods (see details below).
We have done a preliminary inspection of the photometric data of the EBs mainly to find evidence and confirm third components. As mentioned in Sect. 3.2, we found clear signs of third companions in three systems that we have classified as SB3, but also evidence of potential contributions from further components in another 15 systems. Our photometric analysis indicates that there are at least eight systems where there is likely a third companion, which would represent a lower limit of 3% of triple systems in our sample. We provide comments on these and other systems in Appendix C.
@@ 5.3.1 Orbital periods
The B-type Binaries Characterisation programme (BBC; Villaseñor et al., 2021) studied 88 of the binary candidates identified by Dunstall et al. (2015) with 29 FLAMES observations over a baseline of 427 d. Villaseñor et al. (2021) found orbital solutions for 50 SB1 and 14 SB2 systems, and potential solutions for 20 more systems. We can compare the distributions of orbital periods for both populations, over a similar baseline. For this, we have selected nine epochs from the BBC campaign covering a baseline of 60 d, without including the low S/N epochs 10, 12, 13, 17 (see their Table 1 and data reduction section). We computed the LS periodogram for the corresponding RVs for all systems in the BBC sample.
The results can be seen in Fig. 11. Similarly to the periods for the BLOeM sample, we distinguish between those systems whose PLS exceeds the 0.1% FAL (turquoise circles, 29 system), those between the 0.1 and 1% thresholds (orange rectangles, 13 systems), and those with PLS below the 1% threshold (purple diamonds, 45 systems). We found excellent agreement for the first two groups. When limiting the results to a maximum period of 60 d, we found 93% agreement for systems with PLS above the 0.1% threshold, and 90% agreement for those between 0.1–1%. For this reason we use the periods from both these groups in the comparison with the BLOeM sample, resulting in a sample of 34 systems.
In the case of the BLOeM periods, our test with the BBC sample reaffirms our confidence in the periods that met our significance criterion. However, we restrict the periods to those highlighted in Fig. 10, not including those with P FAL. We do include the orbital periods from OGLE, for a total of 57 objects.
The cumulative distribution of orbital periods is shown in Fig. 12. For comparison, we also plotted the distribution of periods for all BBC systems with a period below 60 d (green triangles). When comparing both samples with nine epochs, there is only a marginal difference around 6 d. To assess the significance of this difference, we have computed the Kolmogorov-Smirnov (KS) test, obtaining a KS-statistic of and a p-value of 0.16, which does not reach significance at the 10% level. However, when comparing with the full BBC sample up to periods of 60 d, we do find significant differences (, p=0.0003). It is clear that both samples with nine epochs lack systems with periods larger than d in contrast to the full BBC sample which approaches a uniform distribution in . Judging from Fig. 11, the significance of the periods found with the LS periodogram starts to decrease after 10 d, and from that point up to 60 d we see a majority of systems with P FAL, even though we did find the correct periods for many of them.
@@ 5.3.2 Distribution of
We also examined the distribution of values for the BBC systems with determined periods. Given the larger number of epochs (29) of the BBC programme and the reliable periods, their values are robust. The distribution is shown in Fig. 13, where we have colour-coded the stacked histogram by the binary classifications used by Villaseñor et al. (2021). The top panel shows the values computed from all the epochs with well determined RV measurements that were used in the computation of orbital solutions. In the bottom panel, we plot the distribution obtained by only considering nine epochs as in the case of the BLOeM sample. Comparing the two panels, we see that the distribution is shifted slightly towards lower values when only nine epochs are used, with long-period, eccentric binaries being the most affected. Indeed, the three SB1 systems that fall below 20 km s-1 all have d. Notably, 34 of the BBC SB1 systems have values below 80 km s-1, using a large binary threshold such as this (see discussion in Sect. 6.2) would miss 64% of the SB1 systems. On the other side, a threshold of 20 km s-1 would still skip some of the confirmed long-period binaries in BBC when only using nine epochs, whereas only three systems without orbital solutions remain above the threshold. This evidence also suggests that 20 km s-1 is an appropriate limit to separate binarity from intrinsic variability.
@@ 6.1 Trend with metallicity
As seen in Sect. 4.3, the large multiplicity fraction we have found for the BLOeM B-type stars suggests a trend with metallicity. To test this dependence, we performed a simple linear regression (as the one applied by Sana et al. 2024 to the O-type stars) to the results from Kobulnicky et al. (2014); Dunstall et al. (2015); Banyard et al. (2022), and our result, where we have adopted as uncertainty for the former the mean of the errors of the other studies, and for simplicity, in the case of asymmetric errors, the mean between them. The fit is shown in Fig. 14, with its 68% confidence interval in shaded blue. We found a slope of , which is away from zero.
To illustrate and compare with the case of solar-type stars, we include in the plot a sample of G- and K-type IV/V stars observed by APOGEE and first presented by Badenes et al. (2018). Close binary fraction values as a function of metallicity were obtained from Moe et al. (2019, their figure 11), with the corresponding fit shown in red. We also include three additional values (red diamonds, not included in the linear regression) at [Fe/H]1.0, 0.2, and 0.5. These points represent a weighted moving average from five different solar-type samples, including the Badenes et al. sample, analysed by Moe et al. (2019, see their figure 18). The averaged values are in good agreement with the former measurements, but we only fit the APOGEE sample since it is much larger ( objects) and presents the lowest errors. For the details of the solar-type samples we refer the reader to Moe et al. (2019), but we note that all samples were corrected for incompleteness up to orbital periods of . We have computed the linear fit to the APOGEE data and found a slope of . The two slope values for solar and B-type stars are not significantly different from each other when compared with a t-test (), suggesting a similar trend of the multiplicity fraction with metallicity. Figure 14 also shows the linear fit found by Sana et al. (2024) for the O-type stars (dashed purple line) with its corresponding MW (Sana et al., 2012), LMC (Sana et al., 2013; Almeida et al., 2017), and SMC (Sana et al., 2024) multiplicity fractions. No clear relation was found for the O-type stars with metallicity by Sana et al. (2024)666For a comparison of binary fractions with the O-type stars and the other BLOeM samples see Appendix A.2.
@@
Close massive binaries ( au) are believed to form through disc fragmentation followed by inward orbital migration via circumbinary accretion (Tokovinin & Moe, 2020). Those in the closest orbits ( d) require extremely early fragmentation, coupled with subsequent envelope mass accretion to harden the orbits to such short periods, a process that remains largely independent of metallicity. This would explain why EBs, which predominantly probe these very close binaries, exhibit multiplicity properties that are invariant with metallicity. This is similar to what we see for the more massive O-type stars. In order to form an O-type star, a high accretion rate and large disc surface density are needed, which do not depend on metallicity. This presents a natural explanation for why the multiplicity fraction of O-type stars remains high at different metallicities.
However, for B-type binaries at wider separations ( au), metallicity might have played a more significant role. Metal-poor accretion discs, such as those in the SMC, are less optically thick and can cool more efficiently. This enhanced cooling and reduced opacity facilitate gravitational instabilities, leading to increased disc fragmentation. Therefore, for stars where fragmentation occurs at later times, there is not enough remaining disc mass or accretion to dissipate the orbital energy effectively and decrease separation beyond au. This mechanism would potentially result in a larger multiplicity fraction in early B-type binaries with separations between 0.2 and 10 au at low metallicity.
@@ 6.2 Considering a higher RV threshold
An alternative explanation is that our assumed distributions of orbital parameters are incorrect. Our results are particularly sensitive to the orbital period distribution, for which we have assumed an Öpik law (see Table 3). As discussed in Sect. 5, the fraction of binary system with periods between –10 d seems to be larger for the BLOeM sample than for the BBC sample of binaries. Although our analysis of the orbital period distribution is inconclusive about the significance of the difference in the distribution of both samples according to the KS test, there are other pieces of supporting evidence that are worth examining. First, the fraction of SB2 systems is much larger among the BLOeM B-type binaries in comparison to the 30 Dor counterpart, with 39% of the binaries (41% if we include the SB3 systems) versus only 17% in the BBC sample. There are several factors influencing the detection of SB2 systems, some are intrinsic to the system such as orbital period, mass ratio, and rotational broadening, but external factors like S/N can also play a role. Villaseñor et al. (2021) estimated a mean S/N for the BBC sample of 33. In contrast, the BLOeM sample has a mean S/N of 55, which is significantly higher, probably responding to the higher fraction of giant stars. Moreover, the fraction of SB2s is similar to that of the VFTS O-type binaries (Almeida et al., 2017), which presented similar S/N to our sample. In conclusion, the fraction of SB2 systems, although large, is not impossible to explain.
Eclipsing binaries are also abundant among the BLOeM B-type III/V sample, reaching 27% of the binaries, compared to only 12% found by the BBC programme (Villaseñor et al., 2021). This could also be seen as evidence of an overabundant fraction of short-period systems. However, contrarily to BBC, our BLOeM sample is dominated by giants which are roughly a factor two larger than MS stars. Since the probability of eclipses scales with radius, giant primaries will have a higher probability of being seen as EBs. Only accounting for MS stars, we find that similar fractions of the BBC and BLOeM binaries are MS EBs (14% and 11% respectively).
@@ 6.3 Binary fraction from the distribution of
We adopt an MC method similar to that described in Sect. 4.3, sampling the same physical and orbital parameters but extending the orbital period range down to to include sub-day periods observed by OGLE. We also extend the mass-ratio distribution to to probe lower-mass companions. We assumed circular orbits () for systems with orbital periods below 2 d (Moe & Di Stefano, 2017), whereas for longer orbits we sample from 0 to (as given by Badenes et al. (2018)). Values of the distribution indices remain as in Table 3. To reduce stochastic noise, we generate 10 000 stars with a binary fraction given by and use a MCMC approach to fit the observed distribution by varying . We scaled each simulated distribution by the ratio of observed to simulated stars, and then compare with the real data using a Poisson-based log-likelihood (Cash, 1979) in log-space intervals:
@@
The posterior distribution for is shown in Fig 15, our simulations clearly favour a high binary fraction of % in good agreement with our value of % reported in Sect. 4.3. Figure 16 shows the observed distribution (black histogram), it can be seen that there is no evident separation between a core and tail component, possibly due to the uncertainty in our RV measurements or the scarcity of very wide binaries () in comparison to the distribution of orbital periods of sun-like stars which peaks roughly at d (see Raghavan et al., 2010; Badenes et al., 2018). We include the distribution of values for four simulated samples of 10 000 stars computed with an intrinsic binary fraction of 0.0, 0.5, 0.79, and 1.0 to illustrate the change in the distribution with . A lower binary fraction naturally increases the number of stars with low-amplitude RV variations, whereas a higher binary fraction adds more highly RV-variable systems. Notably, the transition value between single- and binary-dominated bins in the simulations occurs near 20 km s-1, supporting the choice of this value as a sensible threshold.
Another point worth mentioning is the inclusion of slightly evolved stars in our sample. As shown in Sect. 4 and Fig. 4, for typical and detached systems with evolved primaries () can exhibit peak-to-peak RVs about 100 km s-1 lower than those with . These findings illustrate how stellar evolution can effectively shift some of the high-amplitude binaries towards lower values in the observed distribution of . Because our simulations do not track orbital evolution with mass transfer or stellar ageing, these post-interaction systems might escape detection. Future BLOeM papers will explore these evolutionary effects in detail. In conclusion, our multiplicity fraction can thus be interpreted as a lower limit due to binary evolution likely removing part of the high- tail from the distribution.
It is also interesting to examine the distribution of values for (blue histogram in Fig. 16); there is a fraction of simulated single stars (9%) with values above 20 km s-1. The RV values are simulated by adding a Gaussian noise with a standard deviation corresponding to the uncertainty of the RV value at each epoch of a real, randomly assigned, BLOeM star. One single bad epoch (e.g. large uncertainty due to low S/N or a bad multi-component fit) can produce a large value. Figure 17 shows the distribution of the simulated sample of single stars in the binary criteria diagram. Out of the 10 000 stars, only four pass both binary criteria, meaning only 0.04% false-positive detections. This exercise not only confirms the high multiplicity fraction of early B-type dwarfs and giants in the SMC, but also further validates the choice of the binary criteria for our sample.
6.4 Variations in the orbital period distribution